OpenStudy (anonymous):

Help please??? Write the equation of a hyperbola with vertices at (-4, 0) and (4, 0) and co-vertices (0, 5) and (0, -5).

OpenStudy (anonymous):

no one likes these conic section problems, they are not that hard

OpenStudy (anonymous):

as a matter of fact, you can pretty much do this one with your eyeballs, if it is clear that the center is \((0,0)\)

OpenStudy (anonymous):

should be clear since that is half way between \((-4,0)\) and \((4,0)\) so your equation is going to look like \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] you only need \(a\) and \(b\) which you are given in the question

OpenStudy (anonymous):

are is a 4 and b 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the equation is \[\frac{ x ^{2} }{ 4^{2}}+\frac{ y ^{2} }{ 5^{2}}=1\]

OpenStudy (anonymous):

no,put a minus sign there